# In constructing our theory of cosmology, our inspiration is the double-cone geometry of [2]. In the context of two-dimensional (Jackiw–Teitelboim) gravity, the metric of the double cone is that of AdS2 metric in Rindler (i.e., black hole) coordinates:

### v1

@@ 3.2 Gravitational one-loop determinant
We are also interested in the one-loop determinant of the metric and bulk fields around this saddle point. An important contribution comes from the metric zero modes, as in [2, 6]. There is a relative time shift $T_{\mathrm{rel}}$ between the two boundaries, which is canonically conjugate to the energy $E=\frac{r_{\mathrm{h}}^{2}}{8G}$, i.e., $\quantity{T_{\mathrm{rel}},E}=1$ [23, 24]. As we are in three dimensions, there is also a relative shift $\phi_{\mathrm{rel}}$ of the angular coordinate between the two boundaries, canonically conjugate to the angular momentum (we have set the angular momentum to zero in our discussion for simplicity). Thus, for fixed values of the energy and angular momentum, we actually have a two-parameter family of classical saddles to integrate over. The range of integration in $T_{\mathrm{rel}}$ is fixed by the periodicity $T$ we have chosen for the $t$ coordinate, and the range of integration of $\phi_{\mathrm{rel}}$ is similarly fixed by the $2\pi$ periodicity of $\phi$, so integrating over this family of solutions gives us a factor of the volume of the torus in the one-loop determinant, $2\pi T$. As in the double cone [2], this implies that the contribution from these saddles grows linearly with $T$
$\mkern 1.3mu\overline{\mkern-1.3muY_{\tilde{E},\Delta E}(\tilde{T})Y_{E,\Delta E}(-T)\mkern-1.3mu}\mkern 1.3mu\simeq T$ (3.7)
@@
Because we chose a contour that satisfies the Kontsevich–Segal criterion, the remaining contribution to the one-loop determinant from nonzero modes will be well-behaved and finite. That is, we expect the spectrum of fluctuations around this contour to have negative imaginary parts so that their contribution to the one-loop determinant goes to one at large $T$. In the next section, we consider explicitly the fluctuations of a scalar field around our contours and find that the spectrum is related to the quasinormal modes of the black hole, as in [6].
@@ 4.1.2 Scalar fields over the new contour in two dimensions
As in the double-cone case, we expect that only quasinormal modes are present in the physical spectrum, as the one-loop determinant should be well-behaved by the general Kontsevich–Segal argument. It would be interesting to give a more direct argument for this; the argument based on conjugation from [6] does not seem to extend straightforwardly to our contour.
@@ 5 Discussion and speculations
$\mkern 1.3mu\overline{\mkern-1.3muY_{\tilde{E},\Delta E}(\tilde{T})Y_{E,\Delta E}(-T)\mkern-1.3mu}\mkern 1.3mu\eqsim T$ (5.2)
@@
Finally, using the modular invariance of two-dimensional CFTs one can give different interpretations to our spectral observable. To apply a modular transformation, one needs to consider a more conventional spectral form factor involving the canonical partition function $Z(\beta+\mathrm{i}T)$. There are two subtleties when doing this: first, it was observed in [2] that the double cone is not really a saddle point for the canonical spectral form factor, due to a pressure towards lower energies. Second, for states that are close to the BTZ threshold, it is not clear whether the usual relation between energy and temperature will be valid. However, there is at least one case where it seems legitimate. Let us consider equation (5.2) at large $T$ and fixed $E$, so that $\tilde{E}$ is becoming very large. So we expect that we can freely interpret $Y_{\tilde{E},\Delta E}(\tilde{T})$ in terms of a canonical ensemble partition function. We can then perform a modular transformation to relate this to low-energy states. In this case our observable (5.2) is interpreted as a connected correlation function between black holes and the light spectrum of the theory (i.e., those states with $\Delta\ll\frac{c}{12}$). We can then ask what sort of light states could contribute to this connected correlator. The light spectrum includes empty AdS (i.e., the identity operator) and its Virasoro descendants. This spectrum is fixed entirely by symmetry and therefore should not contribute to a connected correlator. So our observable is interpreted as describing correlations between the spectrum of black hole microstates and light, non-vacuum primary states in the theory.
@@ Acknowledgements
We thank X. Dong, T. Hartman, J. Maldacena, S. Shenker, and especially C. Johnson for useful conversations. The work of J.C. is supported by Simons Collaboration on the Nonperturbative Bootstrap. S.F.R. is supported in part by STFC through grant ST/T000708/1. A.M. and K.N. are supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), funding reference number SAPIN/00047.
@@ Appendix A Contours in higher dimensions
In section 2, we considered contours in the complex $u$-plane in two and three dimensions. It is natural to ask what the analogs are in higher dimensions. It turns out that the previous discussion is quite special to low dimensions. One significant reason is that in the BTZ case, the region (III) in figure 3 involves flipping a minus sign for the metric components $g_{tt}$ and $g_{\phi\phi}$ as compared to region (I). For the AdSd+1 black hole with $d>2$, we have $d>1$ transverse directions, and consequently, the contour in region (III) does not satisfy the KS criterion (2.5).
@@
The analog of the $u$-coordinate used previously is an isotropic coordinate, covering the whole Einstein–Rosen bridge. With $r=r_{\mathrm{h}}\sqrt{\cosh u}$, we have
@@
The Einstein–Rosen bridge is given by real $u\in(-\infty,\infty)$. The region with $u=\mathrm{i}\xi$ is a cosmology, propagating from a big bang at $\xi=0$ to a big crunch at $\xi=\frac{\pi}{2}$. The novelty here is that the proper length in the $t$ direction blows up at $\xi=\frac{\pi}{2}$; this is associated with the negative power of $r$ in $f(r)$.
As expected, the horizontal contour at $u=\rho+\mathrm{i}\frac{\pi}{2}$ is not a Lorentzian section—it has three timelike directions. Hence, it violates the Kontsevich–Segal criterion and no allowed contour can approach it. In general, we have
@@
Figure 8: The crossover curve (lower curve) and the limit of all allowed contours (upper curve) for AdS5-Schwarzschild.
For small $y$, this is identical to the AdS3 case. However, there is a crossover when the sign in the second term changes. At large $x$, this occurs when $y=\frac{\pi}{3}$. In general, the crossover is at $\tan^{2}y=2+\tanh^{2}x$. Above this curve, we have
[…]
At large $x$, for $y>\frac{\pi}{3}$ the right-hand side is larger than $\pi$ even for $\frac{\mathrm{d}y}{\mathrm{d}x}=0$, so no allowed contour can reach this region. The curve where $\sum_{\mu=1}^{5}\absolutevalue{\text{Arg}\,\lambda_{\mu}}=\pi$ for $\frac{\mathrm{d}y}{\mathrm{d}x}=0$ (which is not a critical contour) is at $\tan^{2}y=2+\coth^{2}x$. No allowed contour can lie anywhere above this curve. These curves are plotted in figure 8.
We could have a curve that starts off close to the real axis at large $x$ and goes up close to $x=0$ following the cosmology, but as $y\to\frac{\pi}{2}$, has to either end at the singularity at $(0,\frac{\pi}{2})$, or cross $x=0$ and return down to go off to $x\to-\infty$ somewhere below the crossover.

### v3

@@ 3.2 Gravitational one-loop determinant
We are also interested in the one-loop determinant of the metric and bulk fields around this saddle point. An important contribution comes from the metric zero modes, as in [2, 6]. There is a relative time shift $T_{\mathrm{rel}}$ between the two boundaries, which is canonically conjugate to the energy $E=\frac{r_{\mathrm{h}}^{2}}{8G}$, i.e., $\quantity{T_{\mathrm{rel}},E}=1$ [23, 24]. As we are in three dimensions, there is also a relative shift $\phi_{\mathrm{rel}}$ of the angular coordinate between the two boundaries, canonically conjugate to the angular momentum (we have set the angular momentum to zero in our discussion for simplicity). Thus, for fixed values of the energy and angular momentum, we actually have a two-parameter family of classical saddles to integrate over. The range of integration in $T_{\mathrm{rel}}$ is fixed by the periodicity $T$ we have chosen for the $t$ coordinate, and the range of integration of $\phi_{\mathrm{rel}}$ is similarly fixed by the $2\pi$ periodicity of $\phi$. More precisely, we also integrate over a window of energies and angular momenta, so the path integral includes a zero mode integral over
$\int_{0}^{\Delta E}\mathrm{d}E\,\int_{0}^{T}\frac{\mathrm{d}T_{\mathrm{rel}}}{2\pi}\qquad\text{and}\quad\int_{0}^{\Delta J}\mathrm{d}J\,\int_{0}^{2\pi}\mathrm{d}\phi_{\mathrm{rel}}.$ (3.7)
The first integral gives a contribution of the form $T\Delta E$, and the second integral gives a contribution of the form $2\pi\Delta J$. While we write the first integral as $T\Delta E$, the expression is actually symmetric between the two boundaries: while the periods $T$ and $\tilde{T}$ in the two asymptotic regions are not equal, the relation between the energies implies that $T\Delta E=\tilde{T}\Delta\tilde{E}$. Hence, integrating over this family of solutions gives us a factor of the volume of the torus in the one-loop determinant, $2\pi T$. As in the double cone [2], this implies that the contribution from these saddles grows linearly with $T$
$\mkern 1.3mu\overline{\mkern-1.3muY_{\tilde{E},\Delta\tilde{E}}(\tilde{T})Y_{E,\Delta E}(-T)\mkern-1.3mu}\mkern 1.3mu\simeq T$ (3.8)
@@
Because we chose a contour that satisfies the Kontsevich-Segal criterion, the remaining contribution to the one-loop determinant from nonzero modes will be well-behaved and finite. That is, we expect the spectrum of fluctuations around this contour to have negative imaginary parts so that their contribution to the one-loop determinant goes to one at large $T$. In the next section, we consider explicitly the fluctuations of a scalar field around our contours and find that the spectrum is related to the quasinormal modes of the black hole, as in [6].
@@ 4.1.2 Scalar fields over the new contour in two dimensions
As in the double-cone case, we expect that only quasinormal modes are present in the physical spectrum, as the one-loop determinant should be well-behaved by the general Kontsevich-Segal argument. It would be interesting to give a more direct argument for this; the argument based on conjugation from [6] does not seem to extend straightforwardly to our contour.
@@ 5 Discussion and speculations
$\mkern 1.3mu\overline{\mkern-1.3muY_{\tilde{E},\Delta\tilde{E}}(\tilde{T})Y_{E,\Delta E}(-T)\mkern-1.3mu}\mkern 1.3mu\simeq T$ (5.2)
@@
Finally, using the modular invariance of two-dimensional CFTs one can give different interpretations to our spectral observable. To apply a modular transformation, one needs to consider a more conventional spectral form factor involving the canonical partition function $Z(\beta+\mathrm{i}T)$. There are two subtleties when doing this: first, it was observed in [2] that the double cone is not really a saddle point for the canonical spectral form factor, due to a pressure towards lower energies. Second, for states that are close to the BTZ threshold, it is not clear whether the usual relation between energy and temperature will be valid. However, there is at least one case where it seems legitimate. Let us consider equation (5.2) at large $T$ and fixed $E$ so that $\tilde{E}$ is becoming very large. So we expect that we can freely interpret $Y_{\tilde{E},\Delta\tilde{E}}(\tilde{T})$ in terms of a canonical ensemble partition function. We can then perform a modular transformation to relate this to low-energy states. In this case our observable (5.2) is interpreted as a connected correlation function between black holes and the light spectrum of the theory (i.e., those states with $\Delta\ll\frac{c}{12}$). We can then ask what sort of light states could contribute to this connected correlator. The light spectrum includes empty AdS (i.e., the identity operator) and its Virasoro descendants. This spectrum is fixed entirely by symmetry and therefore should not contribute to a connected correlator. So our observable is interpreted as describing correlations between the spectrum of black hole microstates and light, non-vacuum primary states in the theory.
@@ Acknowledgements
We thank X. Dong, T. Hartman, J. Maldacena, S. Shenker, and especially C. Johnson for useful conversations, and M. Sasieta and O. Janssen for comments on previous versions of the paper. The work of J.C. is supported by Simons Collaboration on the Nonperturbative Bootstrap. S.F.R. is supported in part by STFC through grant ST/T000708/1. A.M. and K.N. are supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC), funding reference number SAPIN/00047.
@@ Appendix A Contours in higher dimensions
In section 2, we considered contours in the complex $u$-plane in two and three dimensions. It is natural to ask what the analogs are in higher dimensions. It turns out that the previous discussion is quite special to low dimensions. One significant reason is that in the BTZ case, region (III) in figure 3 involves flipping a minus sign for the metric components $g_{tt}$ and $g_{\phi\phi}$ as compared to region (I). For the AdSd+1 black hole with $d>2$, we have $d>1$ transverse directions, and consequently, the contour in region (III) does not satisfy the KS criterion (2.5).
@@
The analog of the $u$-coordinate used previously is an isotropic coordinate, covering the whole Einstein-Rosen bridge. With $r=r_{\mathrm{h}}\sqrt{\cosh u}$, we have
@@
The Einstein-Rosen bridge is given by real $u\in(-\infty,\infty)$. The region with $u=\mathrm{i}\xi$ is a cosmology, propagating from a big bang at $\xi=0$ to a big crunch at $\xi=\frac{\pi}{2}$. The novelty here is that the proper length in the $t$ direction blows up at $\xi=\frac{\pi}{2}$; this is associated with the negative power of $r$ in $f(r)$.
As expected, the horizontal contour at $u=\rho+\mathrm{i}\frac{\pi}{2}$ is not allowed. The time and sphere directions have a factor of $\mathrm{i}$ here, so $\sum_{\mu=1}^{5}\absolutevalue{\text{Arg}\,\lambda_{\mu}}=2\pi$ along this curve. In general, writing $u=x+\mathrm{i}y$,
@@
Figure 8: The limit of all allowed contours for AdS5-Schwarzschild.
There is a region where the RHS is $>\pi$ even for horizontal contours, where there are no allowed contours. The boundary of this region is at $\tan^{2}y=\coth^{2}x-2$, which is plotted in figure 8. We see that there are no contours that satisfy the KS condition which reach large $x$. The real axis, $u=x$, has $\sum_{\mu=1}^{5}\absolutevalue{\text{Arg}\,\lambda_{\mu}}=\pi$, on the margin of satisfying the KS condition, but for $\coth^{2}x<2$ deforming away from the real axis increases $\sum_{\mu=1}^{5}\absolutevalue{\text{Arg}\,\lambda_{\mu}}$ | (near the origin, for $\coth^{2}x>2$, it decreases it).
This provides a significant obstruction to constructing contours that satisfy the KS condition in higher dimensions, implying also that the double cone contour of [2] does not satisfy this condition in higher dimensions.